OpenStudy (anonymous):

lim t/(radical4t^2+1) t->-infinite evaluate the limit??

OpenStudy (anonymous):

Here's the sloppy version: as t approaches infinity the +1 doesn't matter anymore, so we have t/sqrt(4t^2)=t/2t=1/2, so the limit is 1/2.